Tom and David have 480 marbles between them. Tom arranges them in columns with 6 in each column, and David arranges them with 9 in each column. The total number of columns is 72. Express the above information in the form of an algebraic equation. What is the number of columns of the second boy? Then columns of second boy is ? Number of marbles of the first boy having 6 in each ? Number of marbles of the second boy having 9 in each =?
Total number of marbles?
Algebraic expression is: ?

Answer by scott8148(6628)

let t=Tom so 480-t=David ... (t/6)+((480-t)/9)=72 ... multiplying by 9 and 6 gives 9t+(6*480)-6t=72*9*6

9t+2880-6t=3888 ... 3t=1008 ... t=336 ... 480-t=144

Tom's columns=336/6=56 ... David's columns=144/9=16